OpenStudy (anonymous):

Write the standard form of the line that passes through the point (-2, 4) and is parallel to x - 2y = 6

OpenStudy (welshfella):

change x - 2y = 6 to the form y = mx + b

OpenStudy (amistre64):

one useful property of parallel lines is that they have the same equation, but differ only by a constant

OpenStudy (amistre64):

in other words: ax + by = c and ax + by = k are parallel lines

OpenStudy (anonymous):

y=1/2x−3

OpenStudy (anonymous):

So now what?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

x - 2y = 6 is the stated line x - 2y = k is parallel to it, and it must satisfy the point (-2,4), so what is k when x=-2 and y=4?

OpenStudy (anonymous):

−x+2y=−6 is the line in standard form, but how do I find the line parallel?

OpenStudy (amistre64):

standard form does not have a negative coefficient on the x variable .....

OpenStudy (amistre64):

x - 2y = 6 is the stated line x - 2y = k is parallel to it, and it must satisfy the point (-2,4), so what is k when x=-2 and y=4?

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

So now what? @amistre64

OpenStudy (amistre64):

then let k = -10 and we have the solution x - 2y = -10 is parallel to x - 2y = 6, and it contains the point (-2,4)

OpenStudy (amistre64):

and, its in standard form: ax + by = c, a >=0

OpenStudy (anonymous):

Thanks so much!

OpenStudy (amistre64):

god luck